Help please.I cant solve this word problem, i think i am using the wrong formula.
A 500-pound watermelon is 99% water. As it was sitting in the sun on a hot summer day, some water evaporated and the watermelon ended up 98% water. What is the new weight of the watermelon?
Please and thank you...
: Help please.I cant solve this word problem, i think i am using the wrong formula.
A 500-pound watermelon is 99% water. As it was sitting in the sun on a hot summer day, some water evaporated and the watermelon ended up 98% water. What is the new weight of the watermelon?
Please and thank you...

Answer by stanbon(19018)

A 500-pound watermelon is 99% water. As it was sitting in the sun on a hot summer day, some water evaporated and the watermelon ended up 98% water. What is the new weight of the watermelon?
-----------
Before evaporation:
0.01*500 = 5 lbs not water: 0.99*500 = 495 lbs of water
------------
After evaporation:
Total weight = x lbs
5 lbs not water: 0.98x of water
------------
EQUATION:
5 + 0.98x = x
5 = 0.02x
x = 250 lbs (total weight after evaporation)
=============
Cheers,
Stan H.